OpenStudy (anonymous):

Todd wants to make a cheesecake for a birthday party. The recipe calls for a 9-inch diameter round pan. Todd only has square pans. He has an 8-inch square pan, a 9-inch square pan, & a 10-inch square pan. (8x8, 9x9, 10x10) Which pan comes closest to the area his recipe suggests? Choose one answer. a. 10 inch b. 8 inch c. 9 inch

OpenStudy (anonymous):

area of round pan= pi*r^2=pi*(9/2)^2=pi*81/4 approximately 64 in^2 take the square root of 64 to see which square pan approximates this: sqrt(64in^2)=8in

OpenStudy (anonymous):

Thank You !! Your right !

OpenStudy (anonymous):

yw